### v5

@@ Abstract
This study examines strategic issues in fair rank-minimizing mechanisms, which choose an assignment that minimizes the average rank of object types to which agents are assigned and satisfy a fairness property called equal treatment of equals. As one of these fair mechanisms, the uniform rank-minimizing mechanism is considered. We focus on the case where agents can refuse their assignment and obtain the outside option instead. Without the refusal option, truth-telling is not strategically dominated by any strategies if a fair rank-minimizing mechanism is employed. However, if agents have the option and the uniform rank-minimizing mechanism is employed, then a strategy called an outside option demotion strategy strategically dominates truth-telling. Moreover, we show that adopting this strategy may lead to inefficient assignments. To counter this, we consider a modification of the uniform rank-minimizing mechanism, though it may lead agents to strategically reduce the number of acceptable types.
@@ 1 Introduction
In this study, we discuss the strategic issues of fair rank-minimizing allocation mechanisms mainly when agents have the option to refuse their assignment determined by one of the mechanisms and obtain their outside option instead. There are several agents and types of object, including a null object type, which corresponds to the outside option. In a mechanism, each agent reveals one preference ranking, which may or may not be their true preference ranking. In rank-minimizing mechanisms, assignments are determined as to minimize the total (or average) revealed rank of the objects to which the agents are assigned.
The rank distribution of assignments is a critical consideration for many real-world matching authorities. For instance, Featherstone (2020) notes that Teach For America takes rank distribution into account when making assignment decisions. Consequently, rank-minimizing mechanisms are discussed in recent studies, including Featherstone (2020), Nikzad (2022), Ortega and Klein (2023), Feizi (2024), and Troyan (2024). Several of these studies highlight the advantages of rank-minimizing mechanisms, particularly regarding the efficiency of assignments. By the results of Featherstone (2020) and Feizi (2024), any rank-minimizing mechanism satisfies other efficiency properties, such as ordinal efficiency, which is first introduced by Bogomolnaia and Moulin (2001). Moreover, Ortega and Klein (2023) compare the outcomes of rank-minimizing mechanisms with those of the DA mechanism and the TTC mechanism, showing that rank-minimizing mechanisms improve upon the others in several aspects of efficiency.
In this study, we focus on rank-minimizing mechanisms that satisfy the fairness property of equal treatment of equals, which ensures that agents revealing identical preference rankings are assigned to each object type with the same probability. [^1] Yokote et al. (2019) mention that this fairness property is a classical concept discussed by Aristotle in Book V of Nicomachean Ethics. This fairness property is discussed by Bogomolnaia and Moulin (2001) and many other previous studies on various allocation problems (Thomson 2011). We consider a slightly stronger version of this property. We call a rank-minimizing mechanism satisfying equal treatment of equals a fair rank-minimizing mechanism. To consider these mechanisms, we adopt the probabilistic assignment framework. Among the mechanisms, we first consider a representative one called the uniform rank-minimizing mechanism. Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also mainly consider this mechanism, at least in the literature, this is the most common rank-minimizing mechanism.
We discuss the strategic manipulation incentives of agents when the assignment is determined by a fair rank-minimizing mechanism. Specifically, we consider the case where agents can refuse their assignment, and if they refuse it, then the outside option is assigned to them. First, as a benchmark case, we discuss the situation where agents cannot refuse their assignment. This case is discussed by many previous studies, such as those mentioned above. Featherstone (2020) shows that any rank-minimizing mechanism is not strategy-proof. On the other hand, Troyan (2024) shows that the uniform rank-minimizing mechanism is not obviously manipulable, which is a strategic property introduced by Troyan and Morrill (2020). Moreover, Nikzad (2022) shows that the uniform rank-minimizing mechanism is Bayesian incentive compatible in uniform markets, which he primarily discusses. Related to but logically independent to their results, we show that, in this case, truth-telling is not strategically dominated by any untrue preference under any fair rank-minimizing mechanism. These results imply that in this case, strategic manipulation is a critical problem of rank-minimizing mechanisms.
Next, as the main analysis, we consider the case where agents have the refusal options. [^2] Feigenbaum et al. (2020) and Afacan (2022) state that in some real school choice markets, such as public school markets in Boston, New York City, and Turkey, students actually refuse their assignments. For example, about 10% of the students withdrew from the New York City public high school placement. Featherstone (2020) theoretically considers the situation, but focuses only on the case where each agent has limited information on both the preferences of other agents and the capacities of all object types except the outside option. In contrast, we assume that each agent has complete information on the capacities of all object types. For example, we consider a school choice market where information about all schools, including their capacities, is disclosed to all students. We do not need any assumptions about the information structure of agents regarding the preferences and rationality of other agents, because we focus only on strategic dominance relations.
We consider an outside option demotion strategy, which involves revealing a ranking where the rank of the outside option is placed at the bottom, without changing the order of the truly acceptable object types. We show that, if the uniform rank-minimizing mechanism is employed and agents have the refusal option, then any outside option demotion strategy weakly strategically dominates truth-telling. Moreover, with some additional conditions, some outside option demotion strategy (strictly) strategically dominates truth-telling.
@@
The reason is as follows: First, for some preference profiles of other agents, an agent can successfully avoid the outside option and obtain a truly acceptable object type because assigning them the outside option would significantly increase the sum of ranks. However, this manipulation carries the risk of being assigned an unacceptable object type. Yet, when agents have the refusal option, they can simply reject such assignments and choose the outside option instead. Thus, in this case, the risk associated with adopting this strategy is entirely eliminated.
Moreover, we show that if an agent adopts an outside option demotion strategy and the uniform rank-minimizing mechanism is employed, an inefficient assignment may be realized. This is because, in that case, an unacceptable type may be assigned to the agent who adopts the strategy, and then the agent refuses it, even though the type is desired by other agents. Therefore, in this case, strategic manipulation is a critical problem for rank-minimizing mechanisms.
Featherstone (2020) shows that revealing the full extension, which is one of the outside option demotion strategies, is a weakly better strategy than any other strategies in the case where the agents have little information and the refusal option. The implication of Featherstone (2020) is that rank-efficient mechanisms have strategic issues if the agents have little information and the outside option is somewhat valued by some agents. On the other hand, by our result, this implication is robust for the uniform rank-minimizing mechanism even without the low information assumption. Moreover, we show that revealing the full extension may be strategically dominated by another outside option demotion strategy because, contrary to our model, Featherstone (2020) assumes that the agents have no information on the capacities of object types other than their outside option.
Due to the negative result on the uniform rank-minimizing mechanism introduced above, we modify it so that truth-telling is not strategically dominated by any other strategy. We call this mechanism the modified uniform rank-minimizing mechanism, which is shown to be fair. That is, to ensure that any outside option demotion strategy can fail in some case, in the mechanism, increasing acceptable object types can make the agent less likely to be assigned their acceptable types. However, using the modified uniform rank-minimizing mechanism may create another strategic issue: an agent may have an incentive to reduce the number of truly acceptable object types.
@@ 2 Model
for all $a\in\mathcal{A}$ and all $o\in\mathcal{O}$. Let $\mathcal{X}$ be the set of all possible assignments. Specifically, $y\in\mathcal{X}$ is said to be a deterministic assignment if $y_{ao}$ is either 00 or $1$ for all $\left(a,o\right)\in\mathcal{A}\times\mathcal{O}$. Let $\mathcal{Y}\subseteq\mathcal{X}$ be a set of all possible deterministic assignments.
@@
for all $k=1,2,\cdots,\left|\mathcal{O}\right|$, then agent $a$ with $p_{a}$ is said to weakly prefer (in the sense of first-order stochastic dominance) $x$ to $x^{\prime}$. Moreover, $a$ is said to strictly prefer $x$ to $x^{\prime}$ if $a$ weakly prefers $x$ to $x^{\prime}$ and (1) is satisfied with inequality for some $k=1,2,\cdots,\left|\mathcal{O}\right|-1$.
@@
The efficiency property of rank-minimizing is stronger than several other efficiency properties, such as rank efficiency and ordinal efficiency, analyzed in this literature (see Figure 1 in Feizi (2024)). Among them, not being wasteful is a very weak property.
In the subsequent section, we focus on the mechanisms whose input is $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and output is rank-minimizing for any $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$. Before that we introduce some technical results.
@@
$\bar{k}\left(p_{a}\right)=\min\left\{k\in\left\{1,\cdots,\left|\mathcal{O}\right|\right\}\text{ }\left|\text{ }\sum\limits_{\bar{o}:R(\bar{o},p_{a})\leq k}q_{\bar{o}}\geq\left|\mathcal{A}\right|\right.\right\},$
@@
If $x\in\mathcal{X}$ be an assignment that is not wasteful for $p$, then $x_{ao}=0$ for all $o$ such that $R(o,p_{a})>\bar{k}\left(p_{a}\right)$.
This means that if an assignment is not wasteful for $p$, then $a$ must be assigned to an object type that ranked higher than or equals to $\bar{k}\left(p_{a}\right)$. Typically, under a nonwasteful assignment, the object types that ranked lower than $\varnothing$ in the preference ranking of agent $a$ are never assigned to $a$, because $\varnothing$ is the object with a sufficient number of copies.
@@ 3 Rank-minimizing Mechanism
In this section, we consider the following game. First, all agents simultaneously reveal a preference order in $\mathcal{P}$, which may or may not be their true preference order $r_{a}\in\mathcal{P}$. Let $p_{a}\in\mathcal{P}$ be the preference order revealed by $a\in\mathcal{A}$ and $p=\left(p_{a}\right)_{a\in\mathcal{A}}\in\mathcal{P}^{\left|\mathcal{A}\right|}$. Second, based on $p$, an assignment is determined via a mechanism.
@@
A mechanism employed in the second stage is represented by $f:\mathcal{P}^{\left|\mathcal{A}\right|}\rightarrow\mathcal{X}$. First, we consider an efficiency property on mechanisms. We call $f$ a rank-minimizing mechanism if $f(p)$ is rank-minimizing for all $p$.
Second, we introduce a fairness property. First, we introduce a well-known one. A mechanism $f$ is said to satisfy weak equal treatment of equals (hereafter weak ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}=p_{a^{\prime}}$ implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$. This requires that two agents revealing the same preference must have the same assignment in the ex-ante sense. This property is called just “equal treatment of equals” by many previous studies such as Bogomolnaia and Moulin (2001) and Troyan (2024). In this study, we mainly use a slightly stronger one.
A mechanism $f$ is said to satisfy equal treatment of equals (hereafter ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}^{k^{\prime}}=p_{a^{\prime}}^{k^{\prime}}$ for all $1,\cdots,\bar{k}\left(p_{a}\right)$ implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$ [^3] Han (2024) also considers a similar property, called symmetricity, which is slightly stronger than equal treatment of equals. Unlike in our model, they take into account the priorities of objects and define agents as equal only if they are equal with respect to the priorities of all objects. By Lemma 2, as long as the assignment is not wasteful, any type that is ranked lower than $\bar{k}\left(p_{a}\right)$th in the preference ranking revealed by $a$ will not be assigned to $a$, because there are sufficient copies of types ranked higher than or equal to $\bar{k}\left(p_{a}\right)$th. Thus, if two individuals reveal rankings that are identical for positions at or above $\bar{k}\left(p_{a}\right)$th rank, those rankings are considered essentially the same, and they should receive the same assignment in an ex-ante sense.
Other fairness properties discussed in the literature include weak envy-freeness and the equal division lower bound. However, as shown by Feizi (2024), both of these properties are incompatible with rank-minimizing. Since ETE is weaker than these fairness properties, it is compatible with rank-minimizing and is therefore adopted in this study.
We focus on the efficient and fair mechanisms; that is, we call $f$ a fair rank-minimizing mechanism if $f(p)$ is rank-minimizing for all $p$ and satisfies ETE. The following mechanism is one of them.
@@
be the uniform rank-minimizing mechanism. By (7),
@@
Thus, $RV\left(f^{U}\left(p\right),p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$; that is, $f^{U}\left(p\right)$ must be rank-minimizing for $p$. Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also discuss the uniform rank-minimizing mechanism, this is one of the most common rank-minimizing mechanisms in the literature.
We have the following result.
@@
The uniform rank-minimizing mechanism satisfies ETE.
Therefore, the uniform rank-minimizing mechanism is a fair rank-minimizing one.
@@
For a mechanism $f$ and an agent $a$, (revealing) $p_{a}^{\prime}$ is said to be weakly strategically dominated by (revealing) $p_{a}$ if $a$ weakly prefers $f\left(p_{a},p_{-a}\right)$ to $f\left(p_{a}^{\prime},p_{-a}\right)$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$. Moreover, for $f$ and $a$, (revealing) $p_{a}^{\prime}$ is said to be (strictly) strategically dominated by (revealing) $p_{a}$ if $p_{a}^{\prime}$ is weakly strategically dominated by $p_{a}$ and there exists some $p_{-a}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $f\left(p_{a},p_{-a}^{\prime}\right)$ to $f\left(p_{a}^{\prime},p_{-a}^{\prime}\right)$.
We discuss whether truth-telling ($r_{a}$) can be strategically dominated by revealing an untrue preference $p_{a}\neq r_{a}$. Fix any $r_{a}$ and any $p_{a}\neq r_{a}$. Let $k$ be the smallest integer such that $p_{ak}\neq r_{ak}$. By Lemma 2, if $k>\bar{k}\left(p_{a}\right)$, then $p_{a}$ and $r_{a}$ are essentially equivalent under a fair rank-minimizing mechanism, because any result of a rank-minimizing mechanism is not wasteful. Thus, we focus only on $r_{a}$ and $p_{a}$ satisfying $k\leq\bar{k}\left(p_{a}\right)$ defined above; that is,
@@
is satisfied for that $k$. In other word, we consider the case where the number of object types that are more preferable than the $k-1$th best one in $p_{a}$ is insufficient to distribute to all agents. In this case, for some $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and some fair rank-minimizing mechanism, $a$ is assigned to an object type that is less preferable or equal to the $k$th-best one.
@@
Suppose that a fair rank-minimizing mechanism is employed. For any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ such that revealing $p_{a}$ strategically dominates truth-telling.
@@
$f^{\ast}\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ $=$ $\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}$
otherwise; that is, if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ is neither $\left(\alpha^{\prime},\alpha,\alpha\right)$ nor $\left(\alpha,\alpha^{\prime},\alpha\right)$ nor $\left(\alpha,\alpha,\alpha^{\prime}\right)$. In words, if the first-best type and the second-best type of the three agents are respectively $o_{1}$ and $o_{2}$ and the third-best type of the two of them is $\varnothing$, then the agent revealing $o_{3}$ as the third-best one obtains $o_{1}$ with probability 00.
@@
We assume that $r_{a_{1}}=\alpha^{\prime}$ and this rank-minimizing mechanism $f^{\ast}$ is employed. First, suppose $q_{o_{2}}=1$. In this case, $f^{\ast}$ satisfies ETE and thus this is a fair rank-minimizing mechanism. In this case, $\bar{k}\left(\alpha\right)=\bar{k}\left(\alpha^{\prime}\right)=3$. Then, for example, $r_{a_{1}}=\alpha^{\prime}$ is not strategically dominated by $\alpha$ and any other preference orders, because

### v6

@@ Abstract
This paper investigates the strategic implications of the uniform rank-minimizing mechanism (URM), an assignment rule that selects uniformly from the set of deterministic assignments minimizing the sum of agents’ reported ranks. We focus on settings in which agents may refuse their assignment and instead receive an outside option. Without the refusal option, we show that truth-telling is not strictly dominated under any fair rank-minimizing mechanism; that is, one satisfying equal treatment of equals. However, introducing the refusal option significantly changes strategic incentives: specific manipulations, called outside option demotion strategies, dominate truth-telling under the URM. Moreover, such manipulations can lead to inefficient outcomes, as desirable objects may be refused by misreporting agents and consequently remain unassigned. To address this issue, we propose a modification of the URM that restores undominated truth-telling, although it introduces incentives to underreport acceptable objects. Our results highlight a fundamental trade-off in the design of fair rank-minimizing mechanisms when agents can refuse their assignments.
@@ 1 Introduction
Many assignment problems such as school choice or job assignment require matching agents with heterogeneous objects based on ordinal preferences. In such contexts, policymakers often seek to minimize the sum (or average) of the ranks of assigned objects. For instance, Featherstone (2020) observes that Teach For America explicitly considers rank distribution in its assignment decisions. However, the authority can optimize only the ranks reported by individuals, which may differ from their true preference ranks. These considerations motivate the study of rank-minimizing mechanisms, which determine assignments to minimize the total reported ranks.
A well-known example is the uniform rank-minimizing mechanism (hereafter, URM), which selects uniformly at random from the set of deterministic assignments that minimize the total rank. The URM satisfies equal treatment of equals (Bogomolnaia and Moulin, 2001), a fundamental fairness notion requiring agents who report identical preferences to receive identical (ex-ante) treatment. Mechanisms that are both rank-minimizing and satisfy this fairness condition are referred to as fair rank-minimizing mechanisms, with the URM being their most representative example.
While it is known that rank-minimizing mechanisms are not strategy-proof (Featherstone, 2020), we mainly study whether truth-telling is a dominated strategy. We first consider the benchmark setting without any refusal options, where each agent must accept the assigned object. In this case, we show that truth-telling is not a dominated strategy.
Our main analysis concerns the case in which agents can refuse their assignment and instead receive an outside option. [^1] Feigenbaum et al. (2020) and Afacan (2022) state that in some real school choice markets, such as public school markets in Boston, New York City, and Turkey, students actually refuse their assignments. For example, about 10% of the students withdrew from the New York City public high school placement. In this setting, we show that a class of strategies called outside option demotion strategies, which place the outside option at the bottom of the reported ranking while preserving the order over acceptable objects, can strictly dominate truth-telling under the URM.
@@
The intuition is as follows: for some preference profiles of other agents, an agent who adopts such a strategy may succeed in avoiding the outside option and instead receive a truly acceptable object, because assigning the outside option to that agent would substantially increase the total rank. In the standard setting, this manipulation would be risky, as the agent might be assigned an unacceptable object. However, with the refusal option in place, the agent can simply reject any such assignment and receive the outside option instead. Thus, the strategies become effectively risk-free.
We further show that such strategies may lead to inefficient outcomes: some objects that are desired by agents may be left unassigned if misreporting agents refuse them. In addition, we provide a concrete example in which an outside option demotion strategy forms part of a pure-strategy Nash equilibrium, reinforcing the strategic vulnerability of the URM with the refusal option.
Finally, we propose a modified version of the URM that prevents truth-telling from being strictly dominated and is fair. However, this mechanism introduces new incentives: agents may benefit by promoting the outside option. Our results thus reveal a fundamental trade-off in the design of fair rank-minimizing assignment mechanisms in the presence of refusal options.
Related Literature
There is an extensive literature on fair assignment in probabilistic assignment problems. Prominent studies include Bogomolnaia and Moulin (2001), Budish et al. (2013), and Nesterov (2017), who explore fairness notions such as weak envy-freeness and the equal division lower bound. However, Feizi (2024) shows that these fairness notions are incompatible with rank-minimization. In contrast, ETE, which we adopt in this study, is compatible with rank-minimizing mechanisms and thus suitable for our analysis.
The efficiency of rank-minimizing mechanisms is recently also well studied. Featherstone (2020) and Feizi (2024) show that rank-minimizing mechanisms satisfy stronger efficiency properties than commonly used ones, including ordinal efficiency, originally introduced by Bogomolnaia and Moulin (2001). [^2] Although Featherstone (2020) introduce the concept of rank-minimizing assignments in the economic literature, a broader body of work exists in mathematics and operations research. See Krokhmal and Pardalos (2009) for a comprehensive survey. Ortega and Klein (2023) compare outcomes under rank-minimizing mechanisms with those from the deferred acceptance and top trading cycles mechanisms, showing that the former improve on the latter two in several dimensions of efficiency.
Regarding strategic behavior, much of the existing literature focuses on settings without the refusal option. In this context, Nikzad (2022) establishes Bayesian incentive compatibility for the URM in uniform markets, and Troyan (2024) shows that the URM is not obviously manipulable, in the sense of Troyan and Morrill (2020). Our result in the no-refusal setting is related but logically independent: we show that under any fair rank-minimizing mechanism, truth-telling is not a dominated strategy.
On the other hand, several studies have pointed out strategic issues associated with rank-minimizing mechanisms even without the refusal option. Tasnim et al. (2024) propose strategies for improving individual profit under rank-minimizing mechanisms and show, using both simulated markets and the real matching market in Amsterdam, that applying these strategies can potentially reduce overall matching performance. Bando et al. (2025) also introduce a computationally efficient algorithm for finding profitable manipulations when agents are pessimistic. Our study complements them by theoretically identifying strategic dominance relations in settings where agents can refuse their assignments.
Featherstone (2020) considers settings where agents can refuse their assignment. His analysis assumes that agents possess limited information—specifically, they are unaware of other agents’ preferences and of the capacities of all objects except their own outside option. Our model departs from this assumption: we suppose that agents know the capacities of all objects, as may be the case in a school choice context where this information is publicly available. Moreover, we abstract from assumptions about agents’ beliefs or rationality regarding others, focusing mainly on strategic dominance.
Featherstone (2020) shows that when information is limited as above and agents have the refusal option, revealing the full extension, which is an instance of an ODS, is weakly better than any alternative. His results imply that rank-minimizing mechanisms are vulnerable to manipulation when outside options are valued to some extent. Our findings show that this vulnerability persists even in the absence of information limitation. Moreover, we show that the full extension may itself be strictly dominated by another ODS when agents are informed about object capacities. This highlights a new dimension of strategic risk that does not arise under Featherstone’s setting.
@@ 2 Model
for all $a\in\mathcal{A}$ and all $o\in\mathcal{O}$. Let $\mathcal{X}$ be the set of all possible assignments. Specifically, $y\in\mathcal{X}$ is said to be a deterministic assignment if $y_{ao}$ is either 0 or $1$ for all $\left(a,o\right)\in\mathcal{A}\times\mathcal{O}$. Let $\mathcal{Y}\subseteq\mathcal{X}$ be a set of all possible deterministic assignments.
@@
for all $k=1,2,\cdots,\left|\mathcal{O}\right|$, then agent $a$ with $p_{a}$ is said to weakly prefer (in the sense of first-order stochastic dominance) $x$ to $x^{\prime}$. Moreover, $a$ is said to strictly prefer $x$ to $x^{\prime}$ if $a$ weakly prefers $x$ to $x^{\prime}$ and (1) is satisfied with strict inequality for some $k=1,2,\cdots,\left|\mathcal{O}\right|-1$.
@@
In the subsequent section, we focus on the mechanisms whose input is $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and output is a rank-minimizing assignment for any $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$. Before that we introduce some technical results.
@@
[…]
@@
If an assignment $x\in\mathcal{X}$ is not wasteful for $p$, then $x_{ao}=0$ for all $o$ such that $R(o,p_{a})>\bar{k}\left(p_{a}\right)$.
This means that if an assignment is not wasteful for $p$, then $a$ must be assigned to an object type that ranked higher than or equals to $\bar{k}\left(p_{a}\right)$. Typically, under a nonwasteful assignment, the unacceptable object types in the preference ranking of agent $a$ are never assigned to $a$, because $\varnothing$ is the object with a sufficient number of copies.
@@ 3 Rank-minimizing Mechanism
In this section, we consider the following game. All agents simultaneously reveal a preference order in $\mathcal{P}$, which may or may not be their true preference order $r_{a}\in\mathcal{P}$. Let $p_{a}\in\mathcal{P}$ be the preference order revealed by $a\in\mathcal{A}$ and $p=\left(p_{a}\right)_{a\in\mathcal{A}}\in\mathcal{P}^{\left|\mathcal{A}\right|}$. Then, based on $p$, an assignment is determined via a mechanism.
@@
A mechanism employed in this game is represented by $f:\mathcal{P}^{\left|\mathcal{A}\right|}\rightarrow\mathcal{X}$. First, we consider an efficiency property on mechanisms. We call $f$ a rank-minimizing mechanism if $f(p)$ is rank-minimizing for all $p$. In this study, we mainly focus on the following fair mechanism.
@@
be the uniform rank-minimizing mechanism (URM). By (16),
@@
Thus, $RV\left(f^{U}\left(p\right),p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$; that is, $f^{U}\left(p\right)$ must be rank-minimizing for $p$. Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also discuss the URM, this is one of the most common rank-minimizing mechanisms in the literature.
We show that the URM satisfies fundamental fairness properties. First, we introduce a well-known one. A mechanism $f$ is said to satisfy weak equal treatment of equals (hereafter weak ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}=p_{a^{\prime}}$ implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$. This requires that two agents revealing the same preference must have the same assignment in the ex-ante sense. This property is called just “equal treatment of equals” by many previous studies such as Bogomolnaia and Moulin (2001) and Troyan (2024).
We can show that the URM satisfies a slightly stronger version. A mechanism $f$ is said to satisfy equal treatment of equals (hereafter ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}^{k^{\prime}}=p_{a^{\prime}}^{k^{\prime}}$ for all $1,\cdots,\bar{k}\left(p_{a}\right)$ implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$. [^3] Han (2024) also considers a similar property, called symmetricity, which is slightly stronger than equal treatment of equals. Unlike in our model, they take into account the priorities of objects and define agents as equal only if they are equal with respect to the priorities of all objects. By Lemma 2, as long as the assignment is not wasteful, any type that is ranked lower than $\bar{k}\left(p_{a}\right)$th in the preference ranking revealed by $a$ will not be assigned to $a$, because there are sufficient copies of types ranked higher than or equal to $\bar{k}\left(p_{a}\right)$th. Thus, if two individuals reveal rankings that are identical for positions at or above $\bar{k}\left(p_{a}\right)$th rank, those rankings are considered essentially identical, and they should receive the same assignment in an ex-ante sense.
We say that a mechanism satisfying ETE is a fair mechanism. We have the following result.
@@
The URM satisfies ETE.
Therefore, the URM is a fair rank-minimizing one.
@@
For a mechanism $f$ and an agent $a$, (revealing) $p_{a}^{\prime}\in\mathcal{P}$ is said to be weakly (strategically) dominated by (revealing) $p_{a}\in\mathcal{P}$ if $a$ weakly prefers $f\left(p_{a},p_{-a}\right)$ to $f\left(p_{a}^{\prime},p_{-a}\right)$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$. For $f$ and $a$, (revealing) $p_{a}^{\prime}$ is said to be (strictly) dominated by (revealing) $p_{a}$ if $p_{a}^{\prime}$ is weakly dominated by $p_{a}$ and there exists some $p_{-a}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $f\left(p_{a},p_{-a}^{\prime}\right)$ to $f\left(p_{a}^{\prime},p_{-a}^{\prime}\right)$. Moreover, for $f$ and $a$, (revealing) $p_{a}$ is a (weakly) dominated strategy if there is some $p_{a}^{\prime}\in\mathcal{P}$ that (weakly) dominates $p_{a}$.
Moreover, we say that for $f$, $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ is a Nash equilibrium if there is no $a$ who strictly prefers $f\left(p_{a}^{\prime},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$ for some $p_{a}^{\prime}\in\mathcal{P}$. Therefore, in this context, we define a Nash equilibrium as a profile of strategies such that, for each agent, the resulting assignment is not first-order stochastically dominated by the assignment that results if the agent unilaterally deviated to any other strategy.
We mainly discuss whether truth-telling ($r_{a}$) is a dominated strategy. Fix any $r_{a}$ and any $p_{a}\neq r_{a}$. Let $k$ be the smallest integer such that $p_{ak}\neq r_{ak}$. By Lemma 2, if $k>\bar{k}\left(p_{a}\right)$, then $p_{a}$ and $r_{a}$ are essentially equivalent under a fair rank-minimizing mechanism, because any result of a rank-minimizing mechanism is not wasteful. Thus, we focus only on $r_{a}$ and $p_{a}$ satisfying $k\leq\bar{k}\left(p_{a}\right)$ defined above; that is,
@@
is satisfied for that $k$. In other word, we consider the case where the number of object types that are more preferable than the k1k-1th best one in $p_{a}$ is insufficient to distribute to all agents. In this case, for some $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and some fair rank-minimizing mechanism, $a$ is assigned to an object type that is less preferable or equal to the $k$th-best one.
@@
Suppose that $f$ is a fair rank-minimizing mechanism. For $f$ and $a$ with any $r_{a}\in\mathcal{P}$, truth-telling is not a dominated strategy.
@@
$\left(f^{\ast}\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)\right)_{ao}$ $=$ $\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}$
for all $a\in\mathcal{A}$ and $o\in\mathcal{O}$ otherwise; that is, if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ is neither $\left(\alpha^{\prime},\alpha,\alpha\right)$ nor $\left(\alpha,\alpha^{\prime},\alpha\right)$ nor $\left(\alpha,\alpha,\alpha^{\prime}\right)$. In words, if the first-best type and the second-best type of the three agents are respectively $o_{1}$ and $o_{2}$ and the third-best type of the two of them is $\varnothing$, then the agent revealing $o_{3}$ as the third-best one obtains $o_{1}$ with probability 0.
@@
We assume that $r_{a_{1}}=\alpha^{\prime}$ and this rank-minimizing mechanism $f^{\ast}$ is employed. First, suppose $q_{o_{2}}=1$. In this case, $f^{\ast}$ satisfies ETE and thus this is a fair rank-minimizing mechanism. In this case, $\bar{k}\left(\alpha\right)=\bar{k}\left(\alpha^{\prime}\right)=3$. Then, for example, $r_{a_{1}}=\alpha^{\prime}$ is not dominated by $\alpha$ and any other preference orders, because